A father wants to be accumulated $25000 for his son's college education. The son will admit to a college after 15 years from now. The bank pays 12% interest compounded quarterly, how much money should be deposited now in the bank by the father?
1
Expert's answer
2020-09-27T14:02:13-0400

Compound interest formula is A= P(1+r/n)nt


Where; A= Accumulated/final amount

P= Principal/initial amount

r= interest rate

n= number of times the interest is compounded per period

t= number of periods lapsed


To calculate the initial amount (P); make P the subject of the formula P= A/(1+r/n)nt

Then substitute the figures as follows :

A=$25000

r= 12%

n=4

t= 15


Therefore; P= 25000/(1+0.03)60

P= 4,243.32


The amount the father should deposit is $ 4,243.32